OpenStudy (anonymous):

I need help in math!! Solve 7x − 2y = −3 14x + y = 14 (4, five sevenths) (4, seven fifths) (five sevenths, 4) (seven fifths, 4) ------------- Use the substitution method to solve the following system of equations: 3x + 5y = 3 x + 2y = 0 (1, 1) (6, −3) (6, 1) (1, −3) ----------------------

OpenStudy (anonymous):

@KyanTheDoodle

OpenStudy (haleyelizabeth2017):

For the first one, all you really need to do to start is multiply the first whole equation by 2. Can you do that and show me it when you're done?

OpenStudy (anonymous):

Samantha is going on vacation for the summer and is trying to choose between two different plans. The first plan costs $450 for 3 days at a hotel and 2 days at an amusement park. The second plan offers 5 days at the same hotel and 3 days at the amusement park for $700. The cost of 1 day at the hotel and the amusement park is the same under both plans. How much does a 1-day trip to the amusement park cost? $50 $75 $150 $200 ----------------------------- What is the value of y in the solution to the following system of equations? 5x − 3y = −3 2x − 6y = −6 −1 0 1 2 -------------------------- Miranda bought some apples and oranges at the store for fruit salad. Owen bought three times as many apples and twice as many oranges. Miranda has 5 fruits in her bag, but Owen has 12 fruits in his bag. How many apples did Miranda purchase? 1 2 3 5 ---------------------------------- Gina, Sam, and Robby all rented movies from the same video store. They each rented some dramas, comedies, and documentaries. Gina rented 11 movies total. Sam rented twice as many dramas, three times as many comedies, and twice as many documentaries as Gina. He rented 27 movies total. If Robby rented 19 movies total with the same number of dramas, twice as many comedies, and twice as many documentaries as Gina, how many movies of each type did Gina rent? 3 dramas, 5 comedies, and 3 documentaries 2 dramas, 6 comedies, and 3 documentaries 1 dramas, 4 comedies, and 6 documentaries 4 dramas, 3 comedies, and 4 documentaries

OpenStudy (haleyelizabeth2017):

...is this a test?

OpenStudy (anonymous):

Umm.... A quiz..

OpenStudy (haleyelizabeth2017):

I'm afraid we can't help you then, I'm sorry. :(

OpenStudy (anonymous):

Ok!!